73. A car passes a point at 1 pm and continues along at a constant speed of 64kilometers per hour. A second car passes teh point at 2 pm and goes on a constant speed of 88 kilometers per hour, followint the same route as the first car. At what time will the second car catch up with the first car?
83.The perimeter of a rectangle is 160 meters. If a new rectangle is formed by halving the length of each of on pair of opposite sides and increasing the remaing pair sides by 30 meters each, then the new rectangle aslso has a perimeter of 160 meters. What were the dimensions of teh original rectangle?

Answer by stanbon(48574)

A car passes a point at 1 pm and continues along at a constant speed of 64 kilometers per hour.
Rate = 64 km/h ; distance = x km ; time = x/64 hrs.
--------------------------------------------------------
A second car passes the point at 2 pm and goes on a constant speed of 88 kilometers per hour, following the same route as the first car.
Rate = 88 km/h ; distance = x km ; time = x/88 hrs
---------------------------------------------------------
At what time will the second car catch up with the first car?
------------
time for 1st car - time for 2nd car = 1 hr.
x/64 - x/88 = 1
88x - 64x = 64*88
24x = 64*88
x = 234 2/3 km
-------------------
1st car time = x/64 = [234 2/3]/64 = 3 2/3 hr.
---------
Final Ans: 1pm + 3 2/3 hr = 4 2/3 hr = 4:40 pm
=============================================

83.The perimeter of a rectangle is 160 meters.
2L + 2W = 160 meters
L + W = 80 meters
--------------------------
If a new rectangle is formed by halving the length of each of one pair of opposite sides and increasing the remaing pair sides by 30 meters each, then the new rectangle also has a perimeter of 160 meters.
2(L/2) + 2(W+30) = 160
L + 2W = 100
------------------------
You have two equations:
L + W = 80
L + 2W = 100
---------------
Subtract 1st from 2nd to solve for W:
W = 20 meters
-------
Since L + W = 80, L = 60 meters.
====================================
What were the dimensions of the original rectangle?
W = 20 meters
L = 60 meters
===================
Cheers,
Stan H.